# Frostman random variables, entropy inequalities, and applications

### v3

@@ 4.3 The general forms
When $\phi(x,y)$ is diagonal, the conclusion can be deduced from Lemma 4.1. In the following, we concentrate on the non-diagonal case. By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Applying Theorem 1.3”, there is some positive constant $\epsilon_{4}=\epsilon_{4}(\phi,s)$ such that
$\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\geq n(s+\epsilon_{4})+O_{s}(\log_{+}C_{1}C_{2})+O(\log_{+}c)+O_{\phi,s}(1).$
Combining (13), one has
$H_{n}(X,Y)\leq H_{n}(X+Y)+H_{n}(X)+O(1)\leq 2H_{n}(X+Y)+O(1).$
@@
In the conclusions of Theorems 1.2’-1.4’, the implied constants in the terms $O_{s}(\log_{+}C)$ or $O_{s}(\log_{+}C_{1}C_{2})$ can be independent of $s$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the range of random variables are bounded away from 0 by some absolute constant. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@
The situation is more subtle for this case. It is hard to expect Theorem 2.38 holds true when the dimension gets smaller. We will use the upcoming theorem as a substitution. The core of the proof will follow arguments from [24] by Shmerkin and Wang. However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge. Recall the definition of $\varphi$ from (7).
@@
The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same line as in the case $s>1/2$. So we omit the details. To prove Theorem 5.1, we need the following variant for measures. For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
@@
Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $0<s\leq 1/2$ and $C\geq 1$. Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq C_{d}C^{-1/s}$ with $C_{d}>0$. Then for any $\zeta>0$, we have that
$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,C_{d}}(\log C)$
@@
We emphasize that the term $O_{s,\zeta,C_{d}}(\log C)$ can not be replaced by $O_{s,\zeta,C,C_{d}}(1)$, because in the proof of the main theorem, we are working with infinitely many random variables satisfying (51), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$. So only an error of the form $\log C$ can be turned into entropy-terms as in (54). The main difficulty in this section, which prevents us from applying results in [24] directly, is due to this reason.
@@
Let $0<s<1$ and $C\geq 1$. Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subset[0,1]$. For any $0<\gamma<1/3$, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
@@
Since $X,Y,X^{\prime},Y^{\prime}$ are independent, one can verify that $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
@@ 5.2.1 Material from Shmerkin and Wang
Lemma 5.4.
[Part of Lemma 2.3 [24]] Let $(a_{m})_{m=1}^{\infty}$ be a sequence of positive numbers such that $\sum_{m=1}^{\infty}a_{m}<\infty$. Suppose $\mu\in\mathcal{P}({\mathbb{R}}^{d})$ is $(2^{-m},s,a_{m})$-robust for $m\in[m_{0},m_{1}]$, where $m_{0}<m_{1}\in\mathbb{N}\cup\{+\infty\}$. If $\epsilon:=\sum_{m=m_{0}}^{m_{1}}a_{m}<1$, then there is a set $X$ with $\mu(X)>1-\epsilon$ such that
@@
[Variant of Theorem 4.8 in [24]] Let $1\leq k\leq d,$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function. Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$. Let $(Y,\nu)$ be a compact measure space, $U\subset\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
@@
is $(D;C)$-adapted at all scales $\tilde{\delta}\in(0,\delta_{1})$. We also assume that $\{(x,y):x\in X^{\prime},y\in Y_{x}\}$ is compact. Fix $\tau,\epsilon>0$. There are a set $Y_{0}\subset Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subset X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
@@
By definition, for a fixed $x$, we have that $(\ker D(\pi_{x})(y))^{\perp}=\pi_{x}(y)^{\perp}$ which implies $V_{x}(y)=\pi_{x}(y)^{\perp}$. Using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of length at most $O(C_{\mu}R)$ (see the following figure).
@@
This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta^{-1}}$, provided that $\eta$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
@@
Note that although this does depend on $c$, we choose $C_{1}$ large enough that it will not matter. This implies that $(\rho_{y})_{X_{y}^{\prime}}$ is $(s+\eta,\delta_{1}^{-1})$-Frostman, so the measure $(\mu_{1}\circ P_{v})_{X^{\prime}_{y}}$ is $(s+\eta,O(\delta_{1}^{-1}))$-Frostman for any vector $v\in S^{1}$ because $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim C_{\mu}^{-1}$, so the intersection of any tube of width $r$ going through $x\in\operatorname{spt}\mu_{1}$ can be covered by $C_{\mu}$ disjoint arcs of length $r$ of the unit circle centered at $x$ (see the following figure).
@@ 5.2.2 Modifications to their argument
Using Lemma 5.12 with $C_{\mu}=C_{d}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
@@
Specifically, similar to the proof of Lemma 5.12, using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gtrsim 1$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of width at most $O(R).$ So by the definition of strong thin tubes, we obtain that there is a set $X$ with $(\mu_{1}\times\nu)(X)>1-(\zeta/M)^{2}$ such that for each $x\in X$, there is a set $Y_{x}$ that has $(\mu_{2}\times\nu)(Y_{x})>1-(\zeta/M)^{2}$ and
@@
This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta_{1}^{-1}}$, provided that $\eta_{1}$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
@@
The analogous statement we get is that there is a $c^{\prime}<2(\zeta/M)$, a set $Y$ and $X=\{(y,x):y\in Y,x\in X_{y}\subset\mathbb{R}^{2}\}$ such that $(\mu_{2}\times\nu)(Y)\geq 1-c^{\prime}$, $(\mu_{1}\times\nu)(X_{y})>1-c^{\prime}$ for each $y\in Y$, the set $X$ is compact and most importantly:
@@
Define the map $f(x,y,z,t)=\sqrt{(x-z)^{2}+(y-t)^{2}}$ and the measure $\kappa$ supported in $\mathbb{R}^{4}$ as $\mu_{1}\times\nu\times\mu_{2}\times\nu$. Let the set $X^{\prime}\subset\mathbb{R}^{4}$ be $X^{\prime}:=\{(y,x):y\in Y,x\in X_{y},d(x,y)\in X^{\prime}_{y}\}$ we have that
@@
Write $\phi(x,y)=\alpha x^{2}+\beta y^{2}$. Let $\epsilon=\epsilon(s)$ to be determined later.
Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in the $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Then
$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta}))).\end{split}$ (66)
@@
$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\delta\ll cr_{2}^{s}\delta^{1-s}.$
Calculation reveals that
@@
By Theorem 1.2’, there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1).\end{split}$ (69)
@@
$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{s}(\log_{+}C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1))}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{s}(\epsilon))}.$ (72)
Now we can take $\epsilon=\epsilon(s)>0$ to be sufficiently small so that $\epsilon^{\prime}-O_{s}(\epsilon)\geq\epsilon$ on the right-hand side of (72). Then
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$ (73)
@@
$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
whenever $\delta$ is sufficiently small depending on $\phi,s$ and $C$. The proof is completed. ∎
@@
$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2}.$
And
$P(X\in E,Y\in B(y,r))\ll_{\phi}\begin{cases}\mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(X\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
@@
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{s}(\log C)+O_{s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’. Moreover, we substitute
@@
$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}.$
@@
$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-(s+\epsilon)}$
@@
$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1,\end{cases}$
@@
As a replacement of (69), Theorem 1.4’ shows that
@@
$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{s}(1)}\delta^{-\epsilon}(\#A)$
@@
In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if either of the following two situations happens: (1) we are only considering the case $s\gg 1$; (2) the condition $A\subseteq[0,1]$ is replaced by $A\subseteq[1,2]$.

### v4

@@ 4.4 More general polynomials
By Lemma 2.30, the random variable $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Theorem 1.3” provide us with some positive constant $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
Combining (8) and (9), one has
$H_{n}(X,Y)\leq H_{n}(X)+H_{n}(X+Y)+O(1)\leq 2H_{n}(X+Y)+O(1).$
@@
In the conclusions of Lemma 4.8 and Theorems 1.2’-1.4’, the implied constants in the terms $O_{d,s}(\log_{+}C)$, $O_{d,s}(\log_{+}C_{1}C_{2})$ or $O_{d,s}(\log_{+}c)$ can be independent of $s$ if we are only considering the case $s\gg 1$. We preserve the dependence on $s$ in these terms for the uniformity with the case $0<s\leq 1/2$.
@@
The situation is more subtle for this case. It is hard to expect Theorem 2.38 holds true when the dimension gets smaller. We will use the upcoming theorem as a substitution. The core of the proof will follow the arguments from [24] by Shmerkin and Wang. However, there are enough differences that we cannot merely apply a single theorem of theirs as a black box, so to address this case, we will introduce various notions and results from their paper, and explicitly mention where our arguments diverge.
@@
Here
$\varphi(u)=\frac{u}{2}+\frac{\sqrt{4+u^{2}}-2}{2},\quad(0<u\leq 1).$
@@
The rest of the proof of Theorem 1.4 for $s\leq 1/2$ will follow the same lines as in the case $s>1/2$. So we omit the details. To prove Theorem 5.1, we need the following variant for measures. For two measures $\mu^{\prime},\mu^{\prime\prime}$ whose supports are bounded regions on ${\mathbb{R}}^{2}$, we denote by $\Delta(\mu^{\prime},\mu^{\prime\prime})$ the distance measure between $\mu^{\prime}$ and $\mu^{\prime\prime}$, i.e.,
@@
Let $\mu_{1}$, $\mu_{2}$, and $\nu$ be $(s,C)$-Frostman probability measures on $[0,1]$, where $0<s\leq 1/2$ and $C\geq 1$. Assume that $\text{dist}(\operatorname{spt}(\mu_{1}),\operatorname{spt}(\mu_{2}))\geq\hat{C}C^{-1/s}$ with $\hat{C}>0$. Then for any $\zeta>0$, we have that
$H_{n}(\Delta(\mu_{1}\times\nu,\mu_{2}\times\nu))\geq n(\varphi(2s)-\zeta)+O_{s,\zeta,\hat{C}}(\log C)$
@@
We emphasize that the logarithmic dependence on $C$ in the error term is essential. Because in the proof of the main theorem, we are working with infinitely many random variables satisfying (70), where the Frostman constants can be large compared to $n$, and are not uniform in $(x,y)$. Only an error of the form $\log C$ can be turned into entropy-terms as in (71). The main difficulty in this section, which prevents us from applying the results in [24] directly, is due to this reason.
@@
Let $0<s<1$ and $C\geq 1$. Let $\mu$ be an $(s,C)$-Frostman probability measure with $\operatorname{spt}(\mu)\subseteq[0,1]$. For any $0<\gamma<1/3$, we can find a partition $\{U_{1},U_{2},U_{3}\}$ of $[0,1]$ such that the following hold:
@@
Since $X,Y,X^{\prime},Y^{\prime}$ are independent, the random variables $X_{U_{i}},Y,X_{U_{j}}^{\prime},Y^{\prime}$ are also independent, and $Z_{E_{ij}}\equiv|(X_{U_{i}},Y)-(X_{U_{j}}^{\prime},Y^{\prime})|$.
@@ 5.2.1 Material from Shmerkin and Wang
Lemma 5.4 (Part of Lemma 2.3 [24]).
Let $(a_{m})_{m=1}^{\infty}$ be a sequence of positive numbers such that $\sum_{m=1}^{\infty}a_{m}<\infty$. Suppose $\mu\in\mathcal{P}({\mathbb{R}}^{d})$ is $(2^{-m},s,a_{m})$-robust for $m\in[m_{0},m_{1}]$, where $m_{0}<m_{1}\in\mathbb{N}\cup\{+\infty\}$. If $\epsilon:=\sum_{m=m_{0}}^{m_{1}}a_{m}<1$, then there is a set $X$ with $\mu(X)>1-\epsilon$ such that
@@
[Variant of Theorem 4.8 in [24]] Let $1\leq k\leq d,$ be natural numbers, $D:[0,d]\to[0,k]$ be a continuous function. Fix $c\in(0,1)$, $C\geq 1$, $0<\delta_{1}\leq 1$. Let $(Y,\nu)$ be a compact measure space, $U\subseteq\mathbb{R}^{d}$ be a bounded domain, and let $F_{y}:U\rightarrow\mathbb{R}^{k}$ be either the radial projection to $y$ or the distance to $y$.
@@
is $(D;C)$-adapted at all scales $\tilde{\delta}\in(0,\delta_{1})$. We also assume that $\{(x,y):x\in X^{\prime},y\in Y_{x}\}$ is compact. Fix $\tau,\epsilon>0$. There are a set $Y_{0}\subseteq Y$ with $\nu(Y_{0})>\max(1-2c^{1/2},0)$, and for each $y\in Y_{0}$ there is a set $X_{y}\subseteq X^{\prime}$ with $\mu(X_{y})>\max\{1-2c^{1/2},0\}$, such that
@@
By definition, for a fixed $x$, we have that $(\ker D(\pi_{x})(y))^{\perp}=\pi_{x}(y)^{\perp}$ which implies $V_{x}(y)=\pi_{x}(y)^{\perp}$. Using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg C_{\mu}^{-1}$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of length at most $O(C_{\mu}R)$ (see the following figure).
@@
This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta^{-1}}$, provided that $\eta$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
@@
Note that although this does depend on $c$, we choose $C_{1}$ large enough that it will not matter. This implies that $(\rho_{y})_{X_{y}^{\prime}}$ is $(s+\eta,\delta_{1}^{-1})$-Frostman, so the measure $(\mu_{1}\circ P_{v})_{X^{\prime}_{y}}$ is $(s+\eta,O(\delta_{1}^{-1}))$-Frostman for any vector $v\in S^{1}$ because $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg C_{\mu}^{-1}$, so the intersection of any tube of width $r$ going through $x\in\operatorname{spt}\mu_{1}$ can be covered by $C_{\mu}$ disjoint arcs of length $r$ of the unit circle centered at $x$ (see the following figure).
@@ 5.2.2 Modifications to their argument
Using Lemma 5.12 with $C_{\mu}=\hat{C}C^{\frac{1}{s}},$ we obtain $\eta$ and $C^{\prime}$ only depending on $s$ and $\zeta$ such that if $t<(\varphi(2s)-\zeta/M)$, for some constant $M$ only depending on $s$ and $\zeta,$ which we choose later, and for some $K$ also chosen later, if $(\mu_{1}\times\nu,\mu_{2}\times\nu)$ has (t,K,1c,0)(t,K,1-c,0)-strong thin tubes then it also has $(t+\eta,C^{\prime}\min\{c,K\}^{C^{\prime}},1-9c^{1/2},0)$-strong thin tubes.
@@
Specifically, similar to the proof of Lemma 5.12, using the assumption that $\text{dist}(\operatorname{spt}\mu_{1},\operatorname{spt}\mu_{2})\gg 1$ we obtain that $\pi_{x}(B(z,R))\cap\operatorname{spt}\mu_{2}$ will lie inside a tube of width at most $O(R).$ So by the definition of strong thin tubes, we obtain that there is a set $X$ with $(\mu_{1}\times\nu)(X)>1-(\zeta/M)^{2}$ such that for each $x\in X$, there is a set $Y_{x}$ that has $(\mu_{2}\times\nu)(Y_{x})>1-(\zeta/M)^{2}$ and
@@
This is enough to apply Theorem 5.9 to $Y,\nu,X,$ and the associated projections, $\pi,$ with $\delta_{1}=(KC\mu)^{-C^{\prime}\eta_{1}^{-1}}$, provided that $\eta_{1}$ is small enough in terms of $u$ and $\zeta.$ This will provide us a $c^{\prime}<2c^{1/2}$, a set $Y^{\prime},$ and $X=\{(y,X_{y}):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $\mu_{2}(Y^{\prime})\geq 1-c^{\prime}$, and $\mu_{1}(X_{y})>1-c^{\prime}$ for each $y\in Y^{\prime}$, the set $X$ is compact and most importantly:
@@
The analogous statement we get is that there is a $c^{\prime}<2(\zeta/M)$, a set $Y$ and $X=\{(y,x):y\in Y,x\in X_{y}\subseteq\mathbb{R}^{2}\}$ such that $(\mu_{2}\times\nu)(Y)\geq 1-c^{\prime}$, $(\mu_{1}\times\nu)(X_{y})>1-c^{\prime}$ for each $y\in Y$, the set $X$ is compact and most importantly:
@@
Define the map $f(x,y,z,t)=\sqrt{(x-z)^{2}+(y-t)^{2}}$ and the measure $\kappa$ supported in $\mathbb{R}^{4}$ as $\mu_{1}\times\nu\times\mu_{2}\times\nu$. Let the set $X^{\prime}\subseteq\mathbb{R}^{4}$ be $X^{\prime}:=\{(y,x):y\in Y,x\in X_{y},d(x,y)\in X^{\prime}_{y}\}$ we have that
@@
Write $\phi(x,y)=\rho_{1}(x)+\rho_{2}(y)$. Let $\epsilon=\epsilon(s)$ to be determined later.
Note that $(x,y)\mapsto x+y$ and $(x,y)\mapsto\phi(x,y)$ are both $O_{\phi}(1)$-Lipchitz on $0,1]^{2}$. We can choose a small number $\eta=\eta(\phi)$ with $0<\eta<1/10$ such that $A_{\eta\delta}+_{G}A_{\eta\delta}$ and $\phi_{G}(A_{\eta\delta},A_{\eta\delta})$ are contained in $(A+_{G}A)_{\delta/10}$ and $(\phi_{G}(A,A))_{\delta/10}$, respectively, for any $\delta>0$. Then
$\begin{split}\mathcal{E}_{\delta}(A+_{G}A)&\gg\mathcal{E}_{\delta}(A_{\eta\delta}+_{G}A_{\eta\delta}),\\ \mathcal{E}_{\delta}(\phi_{G}(A,A))&\gg\mathcal{E}_{\delta}(\phi_{G}(A_{\eta\delta},A_{\eta\delta})).\end{split}$ (83)
@@
$\mathcal{L}^{1}(B(y,r_{2})\cap A_{\eta\delta})\ll Cr_{2}^{s}(\#A)\cdot\eta\delta\ll_{\phi}Cr_{2}^{s}\delta^{1-s}.$
The other cases can be calculated in a similar way. Computation reveals that
@@
By Theorem 1.2’ (see Section 4.3), there exists an $\epsilon^{\prime}=\epsilon^{\prime}(s)$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$ (86)
@@
$\qquad\gg 2^{n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon^{\prime}-O_{d,s}(\epsilon))}.$ (89)
Now we can take $\epsilon=\epsilon(d,s)>0$ to be sufficiently small so that the term $\epsilon^{\prime}-O_{d,s}(\epsilon)\geq\epsilon$ on the right-hand side of (89). Then
$\max\{\mathcal{E}_{\delta}(A),\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}$ (90)
@@
$\max\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
whenever $\delta$ is sufficiently small depending on $\phi$ and $s$ and $C$. The proof is completed. ∎
@@
$P(X\in E)\gg_{\phi}\mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot\delta^{1-s+\epsilon}\cdot\delta^{2s-2},$
where the second term on the right-hand side is due to (3). Moreover,
$P\big(X\in E,Y\in B(y,r)\big)\ll_{\phi}\begin{cases}\mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot r\cdot\delta^{2s-2-\epsilon},\quad&\text{if }0<r\leq\delta,\\ \mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot Cr^{s}\delta^{1-s}\cdot\delta^{2s-2-\epsilon}&\text{if }\delta<r\leq 1,\\ \mathcal{L}^{1}(E\cap A_{\eta\delta})\cdot\delta^{1-s}\cdot\delta^{2s-2-\epsilon},&\text{if }r>1.\end{cases}$
@@
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon^{\prime})+O_{d,s}(\log C)+O_{d,s}(\epsilon\log(1/\delta))+O_{\phi,s}(1)\end{split}$
for some $\epsilon^{\prime}=\epsilon^{\prime}(\phi,s)>0$, which results from Theorem 1.3’ (see Section 4.4). Moreover, we substitute
@@
$\max\left\{\mathcal{E}_{\delta}(G)^{1/2},\,\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}.$
@@
$\max\left\{\mathcal{E}_{\delta}(A+_{G}A),\,\mathcal{E}_{\delta}(\phi_{G}(A,A))\right\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-(s+\epsilon)}$
@@
$P(X\in B(x,r))\ll_{\phi}\begin{cases}r\cdot\delta^{s-1-\epsilon},\quad&\text{if }0<r\leq\delta,\\ Cr^{s}\delta^{1-s}\cdot\delta^{s-1-\epsilon},\quad&\text{if }\delta<r\leq 1,\\ 1,\quad&\text{if }r>1\end{cases}$
@@
As a replacement of (86), Theorem 1.4’ (see Section 4.4) shows that
@@
$\max\{\mathcal{E}_{\delta}(B+B),\,\mathcal{E}_{\delta}(\phi(B,B))\}\gg_{\phi,s}C^{-O_{d,s}(1)}\delta^{-\epsilon}(\#A)$
@@
In the conclusions of Theorems 1.5-1.7, the term $C^{-O_{s}(1)}$ can be replaced by $C^{-O(1)}$ if we are only considering the situation $s\gg 1$.